# A Robust Autoencoder Ensemble-Based Approach for Anomaly Detection in Text

### v1

@@ 4.1 Experimental protocol
$\mathcal{A}_{p}(\zeta)=\{\mbox{P}(\zeta)\neq\mbox{P}(\mathrm{y})|(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (8)
@@
$\mathcal{A}_{c}(\zeta)=\{\mbox{P}(\zeta)=\mbox{P}(\mathrm{y})|(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y},\mathcal{A}\backslash\mathcal{A}_{p}\}$ (9)
Dataset Topics Hierarchy
20 Newsgroups 20 Yes
DBpedia 14 14 Yes
Reuters-21578 90 No
Web of Science 134 Yes
Enron 2 No
SMS Spam 2 No
IMDB 2 No
SST2 2 No
Table 1: Presentation of datasets from the literature on outlier detection and the inherent tasks. We describe these corpora by indicating the existence of a topic hierarchy in the labels of the original corpus.
@@ 4.2 Data
Although there are dedicated datasets for outlier detection, such as ODDS or UCI, they mainly provide multidimensional data, time series and computer vision data. Applications such as spam detection and text classification have a rich set of corpora available. Recent work Lai et al. (2020); Ruff et al. (2019a); Kannan et al. (2017); Mahapatra et al. (2012) uses classification datasets such as Reuters-21578111 http://www.daviddlewis.com/resources/testcollections/reuters21578/ and 20 Newsgroups222 http://qwone.com/~jason/20Newsgroups/ with a dedicated preparation in order to compare their approaches.
We use the corpus presented in the section and for each available category, we apply the preparation of independent outliers and contextual outliers with TAC (Algorithm 1). To be fair with each method and each dataset, we set the size of the preparation subset to $350$ and the results are averaged over $10$ of runs. The data is pre-processed by removing lower case and stop words. The train part of each corpus is used for training and evaluation. The TF-IDF model is applied to the whole corpus and only tokens that appear at least three times are kept in the vocabulary. In a first step, we set $\nu=0.10$.
@@
RLEA 0.201 0.642 0.420 0.757 0.334 0.738 0.368 0.747
ErLA 0.325 0.718 0.609 0.880 0.687 0.921 0.840 0.951
Table 2: Results of state of the art models for independent and contextual anomalies with the contamination rate $\nu=0.10$. Area under ROC (AUC) is the default evaluation metric for independent contamination, and Average Precision (AP) is also added for contextual contamination. The experimental study is performed on Distill RoBERTA. Each result is performed on test split prepared through Algorithm 1.
20 Newsgroups.
For 20 Newsgroups we separate the subtopics into seven main topics: computer, forsale, motors, politics, religion, science, sports. We do not count the topic "forsale" for contextual outliers because it has no subtopics.
Reuters-21578.
DBpedia 14.
For DBpedia 14 we create the topic hierarchy based on the ontology provided333mappings.dbpedia.org/server/ontology/classes/ and has six parent topics.
Web Of Science.
Web of Science is often used as a reference for hierarchical classification and provides three levels of topic hierarchy. The third level topics are distributed among the corresponding first level parents. Thus, seven parent topics are present and for child topics that are associated with more than one parent, we keep the largest set of children and delete the others.
@@
Figure 2: Results of our experimental study with $\nu=0.1$, split size of $350$ and number of base detector of $25$. The performance metric is AUROC (AC) and the text representation is RoBERTA.
@@
Figure 3: Boxplots of results of our experimental study with $\nu=0.1$, split size of $350$ and number of base detector of $25$. The performance metric is AUPRC (AP) and the text representation is RoBERTA.
4.3 Setup
We use TAC for preparing contextual contamination on each candidate inliers possible with $\nu=0.1$ (TAC1) and a split size of min$(|D|,1000)$. All results are performed on AUROC and AUPRC reference works from the previous section. We integrate results of one-class autoencoder (OC-AE) and we also benchmark results on a randomly connected autoencoder ensemble (RandAE) Chen et al. (2017). The architecture is similar to Chen et al. (2017) and the autoencoders are following their settings. We also keep the number of runs for each corpus and each split contamination to $10$.
Implementation
For ErLA and RandAE we setup similarly than with the autoencoder and we set the number of base predictors to $20$. We provide the code of our approach [^4] https://github.com/jrmip/ErLA/ using the PyOD base implementation Zhao et al. (2019b). We propose an implementation based on the BaseDetector of PyOD so each of the compared model can be accordingly tested. Randomly connected neural networks can be difficult to implment. Thus, we have implemented three version of our randomly connected autoencoder: one with setting the prune connections to 0 before each backpropagation, another one with the usage of PyTorch built-in prune module and finally a version with Torch-Pruning Fang et al. (2023). Based on our experiments, we use the built-in Pytorch module for reproducibility. Additionally, we also set hyperparameters $\lambda_{1}=0.1$, $\lambda_{2}=0.1$ and $\lambda_{3}=0.05$. For avoiding manifold collapse problem and degenerates solutions, we advise that $\lambda_{3}<\lambda_{1}$. We set the latent dimension of the LNE layer to $32$. On the other hand, we set the epoch number to $50$ and random connection probability between $[0.2,0.5]$.
(a) DBpedia 14
(b) 20 Newsgroups
(c) Web of Science
(d) Reuters-21578
Figure 4: Number of base detectors considering AUC performances on contextual contamination.
(a) DBpedia 14
(b) 20 Newsgroups
(c) Web of Science
(d) Reuters-21578
Figure 5: Number of k-nearest neighbours for one RLAE considering AUC performances on contextual contamination and latent dimension of 32.
@@ 4.3 Setup
Baseline
We propose to compare our approach against OCSVMSchölkopf et al. (2001), a simple one-class autoencoder (OC-AE), RSRAELai et al. (2020) and an ensemble autoencoder model RandAE based on Chen et al. (2017). For OCSVM OC-AE we use the implementation from PyODZhao et al. (2019b) and for RSRAE we use our PyTorch implementation of their code. All our results approximate theirs. We rigorously follow the guidelines provided by Lai et al. (2020).
Alternatively, we propose to use a variant of Algorithm 1 considering $P(y)=P(\zeta)$ for independent contamination. Finally we propose to benchmark our results on corpora presented in Table 1.
We also display CVDD Ruff et al. (2019b) results with several adaptations. In this approach, CVDD learn on all the training corpus instead of the training split. Considering this property, we use their implementation knowing the approach has advantage against other baseline models.
4.4 Results
We propose to present our results on three principal points: independent contamination, contextual contamination and robustness of model scores. Table 2 displays results on TAC1 (contamination of 0.1) and independent contamination. We can see that ErLA is outperforming all models. While the results are similar, our approach is notably standing over the others on SMS Spam and Reuters-21578. Thus, our ensemble method presents success for semantic related, spam and sentiment corpora.
Table 2 displays the experimental results conducted with our approach ErLA. We observe that our approach is outperforming others model with AUROC metric and AUPRC metric. We can see that usage of ErLA allow to mitigate unstable decision of the original RSRAE. We can also see significant difference of performance with Web of Science corpus and Reuters-21578. Additionally, we can observe that the original one-class autoencoder highly benefits from randomly connection and ensemble technique, as it close the gap with other models.
While our performances are competitive, the principal purpose of tackling outlier detection with ensemble methods is to mitigate the bias-variance trade-off. We propose to compare the model results with box-plots, similarly to the previous chapter. The main objective of our contribution is to robust outlier scores for contextual outliers with text. The Figure 2 and the Figure 3 displays an outperforming results from our approach. We can see that the variance of our model is noticeable as the box variance are always smaller than its competitors. Also, the min and max possible scores are close from the median scores, concluding to see that our approach is more efficient, more robust and can handle well language model like RoBERTA.
We can note that reference results are considerably different from ours. The principal reason is due the the contamination process for performing experimental study. We can observe that the literature contaminate both independent anomalies and contextual anomalies without distinctive analysis.
5 Discussion
In this section we propose an observation of various properties of our autoencoder RLAE and its impact inside ErLA (ensemble).
@@ 5 Discussion
The Figure 4 displays the performance impact against the base detector number in ErLA. We can see that starting from 30, there is not substantial gains to increase the number of base detector in the ensemble. While it can slowly increase AUC, it considerably increase the computation time. 20 is then a concrete spot.
@@
In Figure 5 we display the impact of the hyperparameter $k$ in RLAE. As stated in the Reuters-21578 results, we can see that this hyperparameter needs to be wisely picked. While a low value seems to be preferable, we can see that our autoencoder is still robust against this hyperparameter.
@@
In Figure 6 is presented sensibility to three hyperparemeter that impact the learning step: number of hidden layer, the $\lambda_{3}$ parameter and the size of latent space in the ErLA model. While variance is tackled in ensemble methods, hidden layers are not a highly sensible parameter considering the computation/performance tradeoff. In Section 4 the hyperparameter $\lambda_{3}<\lambda_{1}$ is set to $0.05$. The curve is the AUC performance against $\lambda_{3}$ for $\lambda_{1}=0.1$. Finally, the learned embedding inside RLAE should be preferably lower than the output dimension of the autoencoder.
@@ 6 Conclusion
In this work we have introduced ErLA, an ensemble approach with robust autoencoders, optimized through LNE for tackling contextual anomaly in text. Conducted experiments demonstrate that our approach outperform all baseline method with less variance.
One perspective is to study the integration of attention head for mitigating the black box problem of our model. It is common, recently, to display text with their corresponding temperature, thanks to recent language model based on transformers.
The representation of text is a key concept that we want to investigate in the near future. Our approach has proven state of the art results and a great robustness against two kinds of outliers and with a small amount of available documents.
Furthermore, we have displayed that reference contributions have not put sufficient effort to the contamination process in their protocol. One promising perspective is to propose an unsupervised approach for generating different kinds of outliers. Also, our work mainly focuses on the semantic structure of text but syntax is also a promising direction.

### v2

@@ 4 Textual Anomaly Contamination (TAC)
$\mathcal{A}_{p}(\zeta)=\{a|\mbox{P}(\zeta)\neq\mbox{P}(\mathrm{y}),(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (8)
@@
$\mathcal{A}_{c}(\zeta)=\{a|\mbox{P}(\zeta)=\mbox{P}(\mathrm{y}),(\mathrm{a,y})\in\mathcal{A}\times\mathrm{Y}\}$ (9)
The Algorithm 1 presents TAC for independent anomalies and considers three hyperparameter: the contamination rate $\nu$, length of the contaminated split $l$ and the inlier topic noted $\zeta$.
5 Experiments
In this section we present several experiments on both independent anomalies and contextual anomalies. We present all corpora and how we perform this study. Results are commented and an ablation study is proposed.
@@ 5.1 Data
Recent work Lai et al. (2020); Ruff et al. (2019); Kannan et al. (2017); Mahapatra et al. (2012) uses classification datasets such as Reuters-21578111 http://www.daviddlewis.com/resources/testcollections/reuters21578/ and 20 Newsgroups222 http://qwone.com/~jason/20Newsgroups/ with a dedicated preparation in order to compare their approaches. For all corpora we apply data contamination of independent anomalies and contextual anomalies with TAC (Algorithm 1). To be fair with each method and each dataset, we set the size of the preparation subset to the maximum number of available document, and results are averaged over $10$ of runs. Data is pre-processed with lowercase and stop words removal. The train split of each corpus is used for training, and test split for evaluation. In a first step, we set $\nu=0.10$.
20 Newsgroups.
For 20 Newsgroups we separate the subtopics into seven main topics: computer, forsale, motors, politics, religion, science, sports.
Reuters-21578.
DBpedia 14.
For DBpedia 14 Zhang et al. (2015) we create the topic hierarchy based on the provided ontology333mappings.dbpedia.org/server/ontology/classes/ and denote six parent topics.
Web Of Science.
Web of Science Kowsari et al. (2017) is often used as a reference for hierarchical classification and provides three levels of topic hierarchy. Thus, seven parent topics are present and for child topics that are associated with more than one parent, we only keep the largest set of children.
Spam.
We use Enron [^4] https://www.cs.cmu.edu/~enron/ dataset and SMS Spam Almeida et al. (2011) dataset. We prepare a train and test split before contaminating ham in spam, and then spam in ham.
Sentiment.
IMDB Maas et al. (2011) and SST2 Socher et al. (2013) are used for benchmarking independent contamination with more difficult scenarios. Preparation of both corpora is similarly performed than for Spam preparation.
5.2 Preparation
We use TAC for preparing contextual contamination on each candidate inliers with $\nu=0.1$. The maximum amount of available documents is retrieved $|D|$. All results are comapred against Area Under Curve ROC (AUC) and Average Precision (AP) metrics. We also keep the number of runs for each corpus and each split contamination to $10$.
Baseline
We propose to compare our approach against OCSVMSchölkopf et al. (2001), RSRAELai et al. (2020), an autoencoder ensemble model RandNet Chen et al. (2017), CVDD Ruff et al. (2019) and FATE Das et al. (2023). For OCSVM we use the implementation from PyODZhao et al. (2019b) and for RSRAE we use our PyTorch implementation of their code. All our results approximate theirs.
We also display CVDD results with several adaptations. In the original paper, CVDD learns on all the training corpus (inlier only) instead of the training split only. Considering this property, we use their implementation knowing the approach has advantage against other baseline models. The same comment goes for FATE, which additionally use few-shot learning and benefits from anomaly awareness.
Implementation
For RoSAE and RandNet we set the number of base predictors to $20$. We provide the code of our approach [^5] https://anonymous.4open.science/r/RoSAE-COLING-B12E using the PyOD base implementation Zhao et al. (2019b). We propose an implementation based on the BaseDetector of PyOD so each of the compared model can be accordingly tested. We use the pruning methods from PyTorch for creating randomly connected autoencoders. Additionally, we also set hyperparameters $\lambda_{1}=0.1$, $\lambda_{2}=0.1$ and $\lambda_{3}=0.05$. For avoiding manifold collapse problem and degenerates solutions, we advise that $\lambda_{3}<\lambda_{1}$. We set the latent dimension of the LNE layer to $32$. On the other hand, we set the epoch number to $50$ and random connection probability between $[0.2,0.5]$.
@@ 5.2 Preparation
FATE 0.301 0.689 0.536 0.867 0.628 0.913 0.784 0.951
RLAE 0.201 0.642 0.420 0.757 0.334 0.738 0.368 0.747
RoSAE 0.325 0.718 0.609 0.880 0.687 0.921 0.840 0.951
Table 1: Independent and contextual contamination results with contamination rate $\nu=0.10$. Area under ROC (AUC) is the default evaluation metric for independent contamination and Average Precision (AP) is also added for contextual contamination. Embedding is Distill RoBERTA and each result is performed on test split through Algorithm 1. FATE uses few-shot learning and has awareness of 5 anomalies.
@@
Figure 3: Experimental study (AUC) with $\nu=0.05$, split size of $350$ and number of base detector of $25$.
@@
Figure 4: Experimental study (AP) with $\nu=0.05$, split size of $350$ and number of base detector of $25$.
5.3 Results
We propose to present our results on three principal points: independent contamination, contextual contamination and robustness of model scores. Table 1 displays results on TAC and independent contamination. We can see that RoSAE succeeds to compete against recent contributions like FATE. While the results are similar, our approach is notably standing over the others on non topic-related corpora. Thus, our ensemble method presents success for semantic related, spam and sentiment corpora on independent contamination.
For contextual contamination we observe that our approach is outperforming others model with AUC metric and AP metric. We can see that usage of RoSAE allows to mitigate unstable decision of the original RLAE. We can also see significant difference of performance with 20 Newsgroups corpus and Reuters-21578. For contextual contamination, our approach outperforms all models. On AP metric, RoSAE presents 9% higher results than top competitor, on average.
(a) Number of base detector
(b) Study of hyperparameter K
Figure 5: Results of RoSAE with contextual contamination on size of ensemble and on different values for hyperparameter K.
@@ 5.3 Results
While our performances are competitive, the principal purpose of tackling anomaly detection with ensemble methods is to mitigate the bias-variance trade-off. For benchmarking the robustness of each method, we use a limited amount of documents in training and a low contamination rate ($\nu=0.05$), presenting a difficult scenario to handle. The Figure 3 and the Figure 4 displays outperforming results from our approach. We can see that the variance of our model is noticeable as the box variance are always similar or smaller than its competitors. Also, the min and max possible scores are close from the median scores, concluding to see that our approach is more efficient and more robust.
Reference results are considerably different from ours. The principal reason is due the the contamination process for performing experimental study. We can observe that the literature contaminate both independent anomalies and contextual anomalies without distinctive analysis. Another observation lies in the competitive results from FATE thanks to its few-shot learning mechanism.
6 Ablation study
In this section we propose an observation of various properties of our autoencoder RLAE and its impact inside RoSAE (ensemble).
@@ 6 Ablation study
The Figure 5(a) displays the performance impact against the base detector number in RoSAE. We can see that starting from 30, there is not substantial gains to increase the number of base detector in the ensemble. While it can slowly increase AUC, it considerably increase the computation time. 20 is then a concrete spot.
@@
In Figure 5(b) we display the impact of the hyperparameter $k$ in RLAE. As stated in the Reuters-21578 results, we can see that this hyperparameter needs to be wisely picked. While a low value seems to be preferable, we can see that our autoencoder is still robust against this hyperparameter for other dataset.
@@
In Figure 6 is presented sensibility to three hyperparemeter that impact the learning step: number of hidden layer, the $\lambda_{3}$ parameter and the size of latent space in the RoSAE model. While variance is tackled in ensemble methods, hidden layers are not a highly sensible parameter considering the computation/performance tradeoff. In Section 5 the hyperparameter $\lambda_{3}<\lambda_{1}$ is set to $0.05$. The curve is the AUC performance against $\lambda_{3}$ for $\lambda_{1}=0.1$. Finally, the learned embedding inside RLAE should be preferably lower than the output dimension of the autoencoder.
@@ 7 Conclusion
In this work we have introduced RoSAE, an ensemble approach with robust autoencoders, optimized through LNE for tackling contextual anomaly in text. Conducted experiments demonstrate that our approach outperform all baseline methods with less variance on contextual contamination. We can also note that results displays that state-of-the-art approaches are performing independent contamination, and present less robust results than RoSAE.
8 Limitations and Ethical Considerations
We admit one major limitation from this work. As we have presented a large number of dataset, we lack of dedicated corpora. While our work is an attempt to reproduce real-world scenario, we only rely to synthetic contamination. On the other hand, anomaly detection is useful for users and experts, but can lead to unfortunate usages in social media.